# Query-Level Uncertainty in Large Language Models

### v3

@@ Abstract
(a) Comparison of performance and running time between our query-level Internal Confidence method and existing answer-level uncertainty measures (Qwen-14B on GSM8K).
@@ 1 Introduction
Clear awareness of knowledge boundaries is central to improving AI, both for efficiency and trustworthiness. The rising usage of LLMs and agents has introduced significant computational and monetary costs (Varoquaux et al., 2025). For example, agentic workflows may cost 5×–25× more per query compared to a simpler LLM prompt (Anthropic, 2025). Regarding efficiency, if LLMs can distinguish known from unknown or simple from hard queries, they can smartly perform adaptive inference to navigate the trade-offs between computational cost and output quality (Chen & Varoquaux, 2024). For queries beyond their parametric knowledge, they can actively trigger RAG to obtain external knowledge (Lewis et al., 2020) or tool calls (Schick et al., 2023). When faced with hard problems, LLMs can engage in slow (or deep) thinking to improve their outputs, which is also known as test-time scaling (Snell et al., 2024; Zhang et al., 2025). Alternatively, they can defer a complex problem to a larger model via model cascading (Dohan et al., 2022; Gupta et al., 2024). This adaptive inference ensures efficient allocation of computational resources, reducing costs while maintaining performance, especially for agentic scenarios. Beyond efficiency, estimating whether a query is answerable also enhances honesty and trustworthiness of LLMs. When faced with highly uncertain queries, models can adopt an abstention strategy (Wen et al., 2024) to withhold potentially misleading responses, important in high-stakes domains like healthcare (Tomani et al., 2024).
Figure 2: Illustrating the difference between answer-level and query-level uncertainty. Query-level uncertainty estimation distinguishes known from unknown queries (knowledge boundary) before generating answers, which is useful for adaptive inference, e.g., efficient RAG, fast–slow reasoning, or cascading models with different abilities.
@@
Prior research has explored different strategies for uncertainty estimation. One line of work learns a probe of internal states to predict uncertainties of queries (Gottesman & Geva, 2024; Kossen et al., 2024). Another branch of work attempts to teach LLMs to explicitly express “I don’t know” in their responses via fine-tuning methods (Amayuelas et al., 2024; Kapoor et al., 2024; Cohen et al., 2024; Zhang et al., 2024a). One common issue of these studies is that they require fine-tuning and training samples, which introduces additional overhead and may restrict their generalizability across models and domains. To address this gap, we introduce a training-free approach to estimate query-level uncertainty that is both simple and effective.
@@
Our approach, termed Internal Confidence, leverages self-evaluation across internal layers and tokens. It is grounded in a simple assumption: LLMs can internally self-assess the boundaries of their knowledge through a single forward pass over the given query, without generating an explicit answer. Inspired by the uncertainty measure $\textrm{P}(\text{True})$ (Kadavath et al., 2022), we prompt LLMs with a yes–no question to self-assess if they are capable of answering a given query, and define the probability assigned to the token Yes as the confidence level, denoted as $\textrm{P}(\text{Yes})$. To fully exploit the latent knowledge within LLMs, our improved Internal Confidence approach computes this sort of $\textrm{P}(\text{Yes})$ at each layer and token position. Subsequently, we aggregate these signals to obtain the overall confidence score. This aggregation is motivated by prior work showing that leveraging logical consistency across layers can improve outputs (Burns et al., 2022; Chuang et al., 2023; Xie et al., 2024). Concretely, we compute a weighted sum across layers and tokens, and the weights are derived from attenuated encoding (Chen et al., 2023), which enables fine-grained control of the influence of adjacent units.
@@
To validate the effectiveness of our proposed Internal Confidence, we conduct experiments on three datasets that cover factual QA and mathematical reasoning tasks. For fair comparison, we adapt existing answer-level methods to the query level. Experimental results demonstrate that our proposed Internal Confidence can distinguish between known and unknown queries more accurately than a range of baselines, while being substantially faster than answer-level approaches (Figure 1a). In terms of applications, we showcase that our proposed method can support efficient RAG and model cascading. On the one hand, Internal Confidence can guide users to assess the trade-offs between cost and quality when invoking additional services. On the other hand, it reveals an “optimal point”, where inference overhead can be reduced without compromising performance (Figure 1b). In conclusion, we introduce the notion of query-level uncertainty and propose a simple yet effective training-free method to estimate it, which enables models to determine whether a query can be addressed without generating any tokens.
@@ 2.1 Uncertainty Estimation and LLMs
Existing approaches to LLM uncertainty primarily focus on estimating the uncertainty of LLM-generated responses, by providing a score intended to reflect the reliability of a query–answer pair (Geng et al., 2024; Shorinwa et al., 2024; Mahaut et al., 2024; Vashurin et al., 2025). These approaches often rely on internal states (Chen et al., 2024a) or textual responses (Kuhn et al., 2023), and commonly use calibration techniques to mitigate issues such as overconfidence (Zhang et al., 2024b) and biases (Chen et al., 2024b). Notably, these methods assess post-generation reliability, i.e., uncertainty regarding a specific answer after it has been produced. In contrast, relatively little work has explored how to quantify a model’s ability to address a query prior to token generation. For example, Gottesman & Geva (2024) propose training a lightweight probe on internal representations to estimate the model’s knowledge about specific entities. Similarly, Semantic Entropy Probes (Kossen et al., 2024) suggest that internal model states can implicitly encode semantic uncertainty, even before any output is generated. To the best of our knowledge, this work is the first to formally define query-level uncertainty and to investigate it systematically.
@@ 3 Problem Statement and Method
In this section, we define the problem and introduce our method, Internal Confidence, a score that reflects whether an LLM can address a query in its own knowledge, prior to generating tokens.
@@ 3.1 Problem Statement
Given a query (including prompt tokens) $\mathbf{x}=(x_{1},\ldots,x_{N})$, we aim to quantify the query-level uncertainty, $U(\mathbf{x})$, without generating an answer $\mathbf{y}$. This differs from existing uncertainty approaches that estimate the uncertainty associated with a specific generated answer, an answer-level uncertainty that can be denoted as $U(\mathbf{x},\mathbf{y})$. We define a query as being within the model’s knowledge boundary if the LLM can produce a correct answer under greedy decoding, i.e., by selecting the highest-probability token at each step without sampling. Conversely, failure to produce the correct answer suggests the query falls beyond the model’s boundary, and it does not possess sufficient knowledge to answer it. While greedy decoding ensures deterministic measurement, it may not always reflect the optimal performance of a model (Song et al., 2024), as alternative decoding strategies like beam search may elicit a better answer. Therefore, this pragmatic framework serves as a heuristic indicator of internal knowledge, rather than an absolute measure. We use this standard to evaluate the estimated query-level uncertainty, i.e., a lower uncertainty indicates a model is more likely to output the correct answer.
Our problem formulation mostly targets epistemic uncertainty of the model, though specific queries and datasets may contain aleatoric effects (see details in Section A). Our study focuses on queries with definite and clear-cut answers, as in factual QA and mathematical reasoning, which have broad applications and allow for clear evaluations. While contentious queries with open and subjective answers are also important in areas such as politics and philosophy, they remain beyond the scope of this work.
@@ 3.2 Method: From $\textrm{P}(\text{Yes})$ to Internal Confidence
Studies have revealed that LLMs can express verbalized uncertainty in their responses (Tian et al., 2023; Xiong et al., 2024), which indicates that LLMs possess an internal mechanism for assessing the correctness of their outputs. Building on this observation, one can explicitly prompt an LLM to self-assess its confidence in answering a given query by constraining the response to a yes–no binary format: “Respond only with ’Yes’ or ’No’ to indicate whether you are capable of answering the {Query} accurately. Answer Yes or No:”. Following that, we can compute the probability assigned to the token $\textrm{P}(\text{Yes})$ at the last token ($x_{N}$):
@@
Here, $N$ is the index of the last token in the query and $L$ is the index of the last layer of the model. $\mathbf{h}_{N}^{(L)}\in\mathbb{R}^{d}$ is the hidden state, where $d$ is the dimensionality of the hidden representations. $\mathbf{W}^{\text{unemb}}\in\mathbb{R}^{|\mathcal{V}|\times d}$ is the unembedding matrix that maps the hidden state $\mathbf{h}_{N}^{(L)}$ to logits over the vocabulary $\mathcal{V}$. The probability $\textrm{P}(\text{Yes})$ can serve as a query-level confidence score here, which is similar to the process of linear probing (Alain & Bengio, 2016), but without any training steps. While this measure is correlated with verbalized uncertainty, a key distinction is that it requires only a single forward pass of the query, without generating any answer tokens.
However, $\textrm{P}(\text{Yes})$ considers only the final hidden state of the LLM, although the intermediate internal states of LLMs preserve rich knowledge and latent information (Chen et al., 2025), especially for uncertainty estimation (Azaria & Mitchell, 2023; Chen et al., 2024a). Furthermore, prior work demonstrates that incorporating logical consistency across layers can improve outputs (Burns et al., 2022; Chuang et al., 2023; Xie et al., 2024).
@@
Figure 3a plots the average $\textrm{P}(\text{Yes})$ of Llama-8B on mathematical queries (the validation set of GSM8K (Cobbe et al., 2021)), across layers and query tokens. [^1] Here, we consider the last $k$ tokens of a query, assuming that a model has seen the entire query and is able to infer its knowledge gap. We observe that the $\textrm{P}(\text{Yes})$ generally increases from lower to higher layers and from left to right positions. If we treat each $\textrm{P}(\text{Yes}\mid\mathbf{h}_{n}^{(l)})$ as a confidence score and compute the Area Under the Curve (AUC), we can obtain an AUC heatmap that illustrates how effectively each internal representation can distinguish known and unknown queries. As shown in Figure 3b, the highest score does not necessarily appear at the top right position. Instead, the representation $\mathbf{h}_{5}^{(27)}$ yields the best AUC, and the performance gradually declines in regions surrounding this point. We refer to this optimal point as the decision center, where the model most effectively separates known from unknown queries.
@@
Figure 3: Left: the internal $\textrm{P}(\text{Yes})$ across tokens and layers. Middle: the AUC of $\textrm{P}(\text{Yes})$ across tokens and layers. Right: decay weights with different localities. Model: Llama-8B; Dataset: GSM8K validation set.
To improve the vanilla $\textrm{P}(\text{Yes})$, we can apply weighted average centering around the decision center, which serves as an ensemble strategy to enhance calibration and expressivity (Zhang et al., 2020; Stickland & Murray, 2020). We refer to this process as Internal Confidence (IC), formally defined as:
@@
To reflect the observation that the AUC performance gradually decays away from the decision center, we adopt Attenuated Encoding, as proposed by Chen et al. (2023), to compute the above weight vectors in Equation 3:
$\delta_{i,j}=\frac{\exp(-\,\alpha\,\left|i-j\right|^{2})}{\sum^{J}_{j=1}\exp(-\,\alpha\,\left|i-j\right|^{2})},$ (4)
where $i$ is the index of the decision center, $\left|i-j\right|$ is the relative distance, and $\alpha>0$ is a scalar parameter that controls the locality value. Locality is a metric that measures the extent to which weights are concentrated in adjacent positions of a center. Given a weight vector $\mathbf{\epsilon}=\{\epsilon_{1},\epsilon_{2},...,\epsilon_{J}\}$ and assuming that the center index is $i$, the locality can be expressed as:
$\text{Loc}(\mathbf{\epsilon})\in\left[0,1\right]=\sum_{j=1}^{J}\frac{\epsilon_{j}}{2^{\left|i-j\right|}}$ (5)
Here, a value of 1 implies that the vector perfectly satisfies the locality property. Figure 3c plots the weights obtained from Equation 4 for varying degrees of locality. This shows that we can account for the influence of neighboring layers and tokens during the averaging process.
@@ 4.1 Settings
For Llama and Qwen, Internal Confidence is computed in the zero-shot setting, whereas for Phi, we use three shots in the prompt, since smaller models benefit from demonstration-based guidance (See details in Section C.2). All LLMs employ greedy decoding to ensure deterministic outputs. The decision center is fixed to the last layer and last token, and we set $\alpha=1.0$ (Equation 4) across all models and datasets.
@@
Method $\uparrow$ AUC $\uparrow$ PRR $\downarrow$ ECE $\uparrow$ AUC $\uparrow$ PRR $\downarrow$ ECE $\uparrow$ AUC $\uparrow$ PRR $\downarrow$ ECE $\uparrow$ AUC $\uparrow$ PRR $\downarrow$ ECE
@@
Figure 5: Impact of locality on validation set performance. We report the average AUC across the three considered datasets. See details in Section C.3.
We elicit responses from the model using a greedy decoding strategy. If the answer aligns with the ground truth, we consider the model as possessing sufficient knowledge and the query as falling within its knowledge boundary. For the first two datasets with short answers, answers are deemed correct if the ROUGE-L (Lin & Och, 2004) of the ground truth is greater than 0.3, which is consistent with prior work (Kuhn et al., 2023). For the GSM8K dataset, we use an LLM evaluator, Mistral-Large (MistralAI, 2024), to assess both reasoning steps and the final answer. Subsequently, each query is paired with a binary label reflecting whether the model is capable of addressing it.
@@
For comparison, we adapt state-of-the-art answer-level methods to quantify the pre-generation uncertainty (see details in Section B): (1) $\text{Max}(-\log p)$ (Manakul et al., 2023), (2) Predictive Entropy (Malinin & Gales, 2021), (3) Min-$K$ Entropy (Shi et al., 2024), (4) Attentional Entropy (Duan et al., 2024), (5) Perplexity, (6) Internal Semantic Similarity (Fomicheva et al., 2020), (7) $\textrm{P}(\text{Yes})$ (top right), corresponding to Equation 1. (8) $\textrm{P}(\text{Yes})$ (naive avg) is a variant of our Internal Confidence that adopts naive averaging to aggregate scores across different tokens and layers.
@@
We evaluate uncertainty by assessing whether a method can distinguish known and unknown queries, which can be treated as ranking problems, i.e., a lower uncertainty means a model is more likely to know the answer to the query. Following prior work (Manakul et al., 2023; Kuhn et al., 2023), we adopt the Area Under the Curve (AUC) and Prediction Rejection Ratio (PRR) (Malinin et al., 2017) as metrics to measure this. Additionally, we compute the Expected Calibration Error (ECE) to assess the calibration of different methods.
4.2 Internal Confidence Can Identify Known and Unknown Queries
Table 1 summarizes the overall results comparing different query-level uncertainty methods. First, we can observe that our proposed Internal Confidence consistently outperforms other baselines in distinguishing known from unknown queries, as reflected in both average AUC and PRR. The advantage becomes more pronounced for larger models such as Llama-8B and Qwen-14B. For instance, on Qwen-14B, it obtains an average AUC of 67.1 and PRR of 31.7, clearly surpassing all other methods. Regarding the calibration (ECE), Internal Confidence is found to consistently achieve a lower error across models and tasks. These findings indicate the effectiveness of Internal Confidence. Finally, we note that the variants, $\textrm{P}(\text{Yes})$ (top right) and $\textrm{P}(\text{Yes})$ (naive avg), generally underperform the full method, which highlights the importance of the attenuated encoding and its decay weights in effectively aggregating signals from different layers and tokens.

### v4

@@ Abstract
(a) Comparison of performance of distinguishing answerable and non-answerable queries and running time between our query-level Internal Confidence method and existing answer-level uncertainty measures (Qwen-14B on GSM8K).
@@ 1 Introduction
Clear awareness of knowledge boundaries is central to improving AI, both for efficiency and trustworthiness. The rising usage of LLMs and agents has introduced significant computational and monetary costs (Varoquaux et al., 2025). For example, agentic workflows may cost 5×–25× more per query compared to a simpler LLM prompt (Anthropic, 2025). Regarding efficiency, if LLMs can distinguish answerable from non-answerable or simple from hard queries, they can smartly perform adaptive inference to navigate the trade-offs between computational cost and output quality (Chen and Varoquaux, 2024). For queries beyond their parametric knowledge, they can actively trigger RAG to obtain external knowledge (Lewis et al., 2020) or tool calls (Schick et al., 2023). When faced with hard problems, LLMs can engage in slow (or deep) thinking to improve their outputs, which is also known as test-time scaling (Snell et al., 2024; Zhang et al., 2025). Alternatively, they can defer a complex problem to a larger model via model cascading (Dohan et al., 2022; Gupta et al., 2024). This adaptive inference ensures efficient allocation of computational resources, reducing costs while maintaining performance, especially for agentic scenarios. Beyond efficiency, estimating whether a query is answerable also enhances honesty and trustworthiness of LLMs. When faced with highly uncertain queries, models can adopt an abstention strategy (Wen et al., 2024) to withhold potentially misleading responses, important in high-stakes domains like healthcare (Tomani et al., 2024).
Figure 2: Illustrating the difference between answer-level and query-level uncertainty. Query-level uncertainty estimation distinguishes answerable from non-answerable queries (knowledge boundary) before generating answers, which is useful for adaptive inference, e.g., efficient RAG, fast–slow reasoning, or cascading models with different abilities.
@@
Prior research has explored different strategies for uncertainty estimation. One line of work learns a probe of internal states to predict uncertainties of queries (Gottesman and Geva, 2024; Kossen et al., 2024). Another branch of work attempts to teach LLMs to explicitly express “I don’t know” in their responses via fine-tuning methods (Amayuelas et al., 2024; Kapoor et al., 2024; Cohen et al., 2024; Zhang et al., 2024a). One common issue of these studies is that they require fine-tuning and training samples, which introduces additional overhead and may restrict their generalizability across models and domains. To address this gap, we introduce a training-free approach to estimate query-level uncertainty that is both simple and effective.
@@
Our training-free and generation-free approach, termed Internal Confidence, leverages self-evaluation across internal layers and tokens. It is grounded in a simple assumption: LLMs can internally self-assess the boundaries of their knowledge through a single forward pass over the given query, without generating an explicit answer. Inspired by the uncertainty measure $\textrm{P}(\text{True})$ (Kadavath et al., 2022), we prompt LLMs with a yes–no question to self-assess if they are capable of answering a given query, and define the probability assigned to the token Yes as the confidence level, denoted as $\textrm{P}(\text{Yes})$. To fully exploit the latent knowledge within LLMs, our improved Internal Confidence approach computes this sort of $\textrm{P}(\text{Yes})$ at each layer and token position. Subsequently, we aggregate these signals to obtain the overall confidence score. This aggregation is motivated by prior work showing that leveraging logical consistency across layers can improve outputs (Burns et al., 2022; Chuang et al., 2023; Xie et al., 2024). Concretely, we compute a weighted sum across layers and tokens, and the weights are derived from attenuated encoding (Chen et al., 2023), which enables fine-grained control of the influence of adjacent units.
@@
To validate the effectiveness of our proposed Internal Confidence, we conduct experiments on three datasets that cover factual QA and mathematical reasoning tasks. For fair comparison, we adapt existing answer-level methods to the query level. Experimental results demonstrate that our proposed Internal Confidence can distinguish between answerable and non-answerable queries more accurately than a range of baselines, while being substantially faster than answer-level approaches (Figure 1a). In terms of applications, we showcase that our proposed method can support efficient RAG and model cascading. On the one hand, Internal Confidence can guide users to assess the trade-offs between cost and quality when invoking additional services. On the other hand, it reveals an “optimal point”, where inference overhead can be reduced without compromising performance (Figure 1b). In conclusion, we introduce the notion of query-level uncertainty and propose a simple yet effective training-free method to estimate it, which enables models to determine whether a query can be addressed without generating any tokens.
@@ 2.1 Uncertainty Estimation and LLMs
Existing approaches to LLM uncertainty primarily focus on estimating the uncertainty of LLM-generated responses, by providing a score intended to reflect the reliability of a query–answer pair (Geng et al., 2024; Shorinwa et al., 2024; Mahaut et al., 2024; Vashurin et al., 2025). These approaches often rely on internal states (Chen et al., 2024a) or textual responses (Kuhn et al., 2023), and commonly use calibration techniques to mitigate issues such as overconfidence (Zhang et al., 2024b) and biases (Chen et al., 2024b). Notably, these methods assess post-generation reliability, i.e., uncertainty regarding a specific answer after it has been produced. In contrast, relatively little work has explored how to quantify a model’s ability to address a query prior to token generation. For example, Gottesman and Geva (2024) propose training a lightweight probe on internal representations to estimate the model’s knowledge about specific entities. Similarly, Semantic Entropy Probes (Kossen et al., 2024) suggest that internal model states can implicitly encode semantic uncertainty, even before any output is generated. To the best of our knowledge, this work is the first to formally define query-level uncertainty and to investigate it systematically.
@@ 3 Problem Statement and Method
In this section, we define the problem and introduce our Internal Confidence, a training-free and generation-free uncertainty that reflects whether an LLM can address a query in its own knowledge, prior to generating tokens.
@@ 3.1 Problem Statement
Given a query (including prompt tokens) $\mathbf{x}=(x_{1},\ldots,x_{N})$, we aim to quantify the query-level uncertainty, $U(\mathbf{x})$, without generating an answer $\mathbf{y}$. This differs from existing uncertainty approaches that estimate the uncertainty associated with a specific generated answer, an answer-level uncertainty that can be denoted as $U(\mathbf{x},\mathbf{y})$. We define a query as being within the model’s knowledge boundary if the LLM can produce a correct answer under greedy decoding, i.e., by selecting the highest-probability token at each step without sampling. Conversely, failure to produce the correct answer suggests the query falls beyond the model’s boundary, and it does not possess sufficient knowledge to answer it. While greedy decoding ensures deterministic measurement, it may not always reflect the optimal performance of a model (Song et al., 2024), as alternative decoding strategies like beam search may elicit a better answer. We stick with greedy decoding for the following reasons: (1) Efficiency – Our method treats a successful greedy decode as a signal that the model knows how to answer, which is a fast proxy. In contrast, non-greedy decoding requires the configuration of beam numbers, probability thresholds, and sampling numbers, which complicates both the definition of knowledge boundary and the assessment cost. (2) Reproducibility – Greedy decoding outputs a single deterministic output for a given input, which offers a stable and reproducible baseline and benchmark.
Therefore, this pragmatic framework serves as a heuristic indicator of internal knowledge, rather than an absolute measure. We use this standard to evaluate the estimated query-level uncertainty, i.e., a lower uncertainty indicates a model is more likely to output the correct answer.
Our problem formulation mostly targets epistemic uncertainty of the model, though specific queries and datasets may contain aleatoric effects (see details in Section A), and the definition of the knowledge boundary is aligned with the parametric knowledge boundary (Li et al., 2024). This boundary of a model is the set of all knowledge encoded in its parameters that can be verified by at least one input–output pair. Our study focuses on queries with definite and clear-cut answers, as in factual QA and mathematical reasoning, which have broad applications and allow for clear evaluations. While contentious queries with open and subjective answers are also important in areas such as politics and philosophy, they remain beyond the scope of this work.
@@ 3.2 Method: From $\textrm{P}(\text{Yes})$ to Internal Confidence
Studies have revealed that LLMs can express verbalized uncertainty in their responses (Tian et al., 2023; Xiong et al., 2024), and they can self-evaluate whether they know the answer to a question without reference to any specific proposed answer (Kadavath et al., 2022), which indicates that LLMs possess an internal mechanism for assessing the correctness of their outputs. At the same time, a recent work indicates that answerable and non-answerable questions are also linearly separable in hidden states (Slobodkin et al., 2023). Building on this observation, one can explicitly prompt an LLM to self-assess its confidence in answering a given query by constraining the response to a yes–no binary format: “Respond only with ’Yes’ or ’No’ to indicate whether you are capable of answering the {Query} accurately. Answer Yes or No:”. Following that, we can compute the probability assigned to the token $\textrm{P}(\text{Yes})$ at the last token ($x_{N}$):
@@
Here, $N$ is the index of the last token in the query and $L$ is the index of the last layer of the model. $\mathbf{h}_{N}^{(L)}\in\mathbb{R}^{d}$ is the hidden state, where $d$ is the dimensionality of the hidden representations. $\mathbf{W}^{\text{unemb}}\in\mathbb{R}^{|\mathcal{V}|\times d}$ is the unembedding matrix that maps the hidden state $\mathbf{h}_{N}^{(L)}$ to logits over the vocabulary $\mathcal{V}$. The unembedding layer provides meaningful and comparable probabilities, whereas the raw logits are not directly interpretable in this way. The probability $\textrm{P}(\text{Yes})$ can serve as a query-level confidence score here, which is similar to the process of linear probing (Alain and Bengio, 2016), but without any training steps. While this measure correlates with verbalized uncertainty, a key distinction is that it requires only a single forward pass of the query, without generating any answer tokens.
However, $\textrm{P}(\text{Yes})$ considers only the final hidden state of the LLM, although the intermediate internal states of LLMs preserve rich knowledge and latent information (Chen et al., 2025), especially for uncertainty estimation (Azaria and Mitchell, 2023; Chen et al., 2024a). Furthermore, prior work demonstrates that incorporating logical consistency across layers can improve outputs (Burns et al., 2022; Chuang et al., 2023; Xie et al., 2024).
@@
Figure 3a plots the average $\textrm{P}(\text{Yes})$ of Llama-8B on mathematical queries—the validation set of GSM8K (Cobbe et al., 2021)—across layers and query tokens. [^1] Here, we consider the last $k$ tokens of a query, assuming that a model has seen the entire query and is able to infer its knowledge gap. We observe that the $\textrm{P}(\text{Yes})$ generally increases from lower to higher layers and from left to right positions. If we treat each $\textrm{P}(\text{Yes}\mid\mathbf{h}_{n}^{(l)})$ as a confidence score and compute the Area Under the Curve (AUROC), we can obtain an AUROC heatmap that illustrates how effectively each internal representation can distinguish answerable and non-answerable queries. As shown in Figure 3b, the highest score does not necessarily appear at the top right position. Instead, the representation $\mathbf{h}_{5}^{(27)}$ yields the best AUROC, and the performance gradually declines in regions surrounding this point. We refer to this optimal point as the decision center, where the model most effectively separates answerable from non-answerable queries.
@@
Figure 3: Left: the internal $\textrm{P}(\text{Yes})$ across tokens and layers. Middle: the AUROC of $\textrm{P}(\text{Yes})$ across tokens and layers. Right: decay weights with different localities. Model: Llama-8B; Dataset: GSM8K validation set.
To improve the vanilla $\textrm{P}(\text{Yes})$, we can apply weighted average centering around the decision center, which serves as an ensemble strategy to enhance calibration and expressivity (Zhang et al., 2020; Stickland and Murray, 2020). We refer to this process as Internal Confidence (IC), formally defined as:
@@
To reflect the observation that the AUROC performance gradually decays away from the decision center, we adopt Attenuated Encoding, as proposed by Chen et al. (2023), to compute the above weight vectors in Equation 3:
$\delta_{j}^{(i)}=\frac{\exp(-\,\alpha\,\left|i-j\right|^{2})}{\sum^{J}_{j=1}\exp(-\,\alpha\,\left|i-j\right|^{2})},$ (4)
where $i$ is the index of the decision center, $\left|i-j\right|$ is the relative distance, and $\alpha>0$ is a scalar parameter that controls the locality value. Locality is a metric that measures the extent to which weights are concentrated in adjacent positions of a center. Given a weight vector $\bm{\delta}^{(i)}=\{\delta_{1}^{(i)},\delta_{2}^{(i)},...,\delta_{J}^{(i)}\}$ and assuming that the center index is $i$, we define its locality as
$\mathrm{Loc}(\bm{\delta}^{(i)})\in[0,1]=\sum_{j=1}^{J}\frac{\delta^{(i)}_{j}}{2^{|i-j|}}.$ (5)
Here, a value of 1 implies that the vector perfectly satisfies the locality property, which means weights are extremely concentrated at the decision center. A low locality means weights are more uniformly assigned to neighborhoods. Figure 3c plots the weights obtained from Equation 4 for varying degrees of locality. This shows that we can account for the influence of neighboring layers and tokens during the averaging process.
@@ 4.1 Settings
For Llama and Qwen, Internal Confidence is computed in the zero-shot setting, whereas for Phi, we use two shots in the prompt, since smaller models benefit from demonstration-based guidance (See details in Section D.2). All LLMs employ greedy decoding to ensure deterministic outputs. The decision center is fixed to the last layer and last token, and we set $\alpha=1.0$ (Equation 4) across all models and datasets.
@@
Method $\uparrow$ AUROC $\uparrow$ PRR $\downarrow$ ECE $\uparrow$ AUROC $\uparrow$ PRR $\downarrow$ ECE $\uparrow$ AUROC $\uparrow$ PRR $\downarrow$ ECE $\uparrow$ AUROC $\uparrow$ PRR $\downarrow$ ECE
@@
We elicit responses from the model using a greedy decoding strategy. If the answer aligns with the ground truth, we consider the model as possessing sufficient knowledge and the query as falling within its knowledge boundary. For the first two datasets with short answers, answers are deemed correct if the ROUGE-L (Lin and Och, 2004) of the ground truth is greater than 0.3, which is consistent with prior work (Kuhn et al., 2023). For the GSM8K dataset, we use an LLM evaluator, Mistral-Large (MistralAI, 2024), to assess both reasoning steps and the final answer. We evaluate the reasoning steps on GSM8K because verifying the reasoning chain is essential to ensure the model truly understands the problem rather than outputting the correct results by chance. Subsequently, each query is paired with a binary label reflecting whether the model is capable of addressing it.
@@
For comparison, we adapt state-of-the-art answer-level methods to quantify the pre-generation uncertainty (see details in Section C): (1) $\text{Max}(-\log p)$ (Manakul et al., 2023), (2) Predictive Entropy (Malinin and Gales, 2021), (3) Min-$K$ Entropy (Shi et al., 2024), (4) Attentional Entropy (Duan et al., 2024), (5) Perplexity, (6) Internal Semantic Similarity (Fomicheva et al., 2020), (7) $\textrm{P}(\text{Yes})$ (top right), corresponding to Equation 1. (8) $\textrm{P}(\text{Yes})$ (naive avg) is a variant of our Internal Confidence that adopts naive averaging to aggregate scores across different tokens and layers.
@@
We evaluate uncertainty by assessing whether a method can distinguish answerable and non-answerable queries, which can be treated as ranking problems, i.e., a lower uncertainty means a model is more likely to know the answer to the query. Following prior work (Manakul et al., 2023; Kuhn et al., 2023), we adopt the Area Under the Receiver Operating Characteristic Curve (AUROC) and Prediction Rejection Ratio (PRR) (Malinin et al., 2017) as metrics to measure this. Additionally, we compute the Expected Calibration Error (ECE) to assess the calibration of different methods.
4.2 Internal Confidence Can Identify Answerable and Non-Answerable Queries
Table 1 summarizes the overall results comparing different query-level uncertainty methods. First, we can observe that our proposed Internal Confidence consistently outperforms other baselines in knowledge boundary detection, as reflected in both average AUROC and PRR. The advantage becomes more pronounced for larger models such as Llama-8B and Qwen-14B. For instance, on Qwen-14B, it obtains an average AUROC of 67.1 and PRR of 31.7, clearly surpassing all other methods. Regarding the calibration (ECE), Internal Confidence is found to consistently achieve a lower error across models and tasks. These findings indicate the effectiveness of Internal Confidence. Finally, we note that the variants, $\textrm{P}(\text{Yes})$ (top right) and $\textrm{P}(\text{Yes})$ (naive avg), generally underperform the full method, which highlights the importance of the attenuated encoding and its decay weights in effectively aggregating signals from different layers and tokens.